There are two consumers on the market: Jim and Donna. Jim’s utility function is U(x, y) = xy. Donna’s utility function is U(x, y) = x 2 y. Income of Jim is mJ = 100 and income of Donna is mD = 150.

a) Find optimal baskets of Jim and Donna when price of y is Py = 1 and price of x is P.
Answer #1

PAGE NO.: DATE: I I 千 Ae ec urubunun m) R. T = mRS P. 2 SO

Optimal bundles:

X=50/P,Y=50

DATE: /I Denwa 2. 2. rm

Optimal bundles:X=100/P,Y=50

Please leave a comment in case of any queries.